### v1

# A parameterized block-splitting preconditioner for indefinite least squares problem

Abstract. We present a stationary iteration based upon a block splitting for a class of indefinite least squares problem. Convergence of the proposed method is investigated and optimal value of the involving parameter is used. The induced preconditioner is applied to accelerate the convergence of the GMRES method for solving the problem. We also analysed the eigenpair distribution of the preconditioned matrix. Some numerical are presented to show the effectiveness of the preconditioner. Numerical comparison with other well-known methods are also presented.
Keywords: Block, splitting, least squares, indefinite, preconditioning, iterative.
AMS Subject Classification
: 65F10, 65F50, 65F08..
1 Introduction
We are concerned with problem of solving the indefinite least squares (ILS) problem
$\min_{x\in\mathbb{R}^{m}}(b-Ax)^{T}J(b-Ax)$ (1)
where $A\in\mathbb{R}^{m\times n}$ with $m\geq n$, and
$Q=\begin{pmatrix}I_{p}&0\\ 0&I_{q}\end{pmatrix},\qquad p+q=m.$ (2)
Problems of this type are encountered in various applications, including the solution of total least squares problems [3, 6, 4] and the field of optimization referred to as Robust $H^{\infty}$ smoothing [4]. If either $p=0$ or $q=0$ the problem is reduced to the classical least squares problem.
The normal equations corresponding to the problem (1) can be written as
$A^{T}JAx=A^{T}Jb.$ (3)
It follows from (3) that unlike the least squares problem, the ILS problem does not necessarily have a solution. The Hessian matrix for the function $f(x)=(b-Ax)^{T}J(b-Ax)$ is $H=A^{T}JA$. So the minimization problem (1) has a unique solution if and only if the matrix $A^{T}JA$ is symmetric positive definite (SPD) (see [4, 7]). When the problem size is small, Chandrasekaran et al. in [4] proposed a direct method based on the QR decomposition and Cholesky factorization for solving the problem. However, when the problem is large, the computational cost becomes prohibitively high. So it is recommended to use iterative methods for solving the problem.
In [9], Liu and Liu presented the two versions of the block SOR method for solving the problem. In [15], Song constructed an unsymmetric SOR (USSOR) method for solving the problem. A variable parameter Uzawa method for solving the indefinite least squares problem was presented by Meng et al. in [12]. Meng et al. in [11] proposed an enhanced Landweber method with momentum acceleration for solving large indefinite least squares problem. Recently, Meng et al. in [10] have stated alternating direction implicit iterative method for the problem.
We assume that the matrix $A$ and the vector $b$ can be written as
$A=\begin{pmatrix}A_{1}\cr A_{2}\end{pmatrix},\qquad A=\begin{pmatrix}b_{1}\cr b_{2}\end{pmatrix}$ (4)
where $A_{1}\in\mathbb{R}^{p\times n}$ is of full column rank, $A_{2}\in\mathbb{R}^{q\times n}$, $b_{1}\in\mathbb{R}^{p}$ and $b_{2}\in\mathbb{R}^{q}$. Xin and Meng in [16] wrote the system (3) in the form
$\mathcal{B}\mathbf{x}=\begin{pmatrix}I&A_{1}&0\cr 0&P&A_{2}^{T}\cr 0&A_{2}&I\end{pmatrix}\begin{pmatrix}\delta_{1}\cr x\cr\delta_{2}\end{pmatrix}=\begin{pmatrix}b_{1}\cr A_{1}^{T}b_{1}\cr b_{2}\end{pmatrix},$ (5)
where $I$ is the identity matrix of appropriate size, $P=A_{1}^{T}A_{1}$ and $\delta=(\delta_{1};\delta_{2})=b-Ax$. Then, they considered the following three block splittings for the above system
$\mathcal{B}$ $=$ $\begin{pmatrix}I&0&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{1}-\mathcal{N}_{1},$ (6)
$\mathcal{B}$ $=$ $\begin{pmatrix}I&0&0\cr 0&P&A_{2}^{T}\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&-A_{1}&0\cr 0&0&0\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{2}-\mathcal{N}_{2},$ (7)
$\mathcal{B}$ $=$ $\begin{pmatrix}I&A_{1}&0\cr 0&P&0\cr 0&0&I\end{pmatrix}-\begin{pmatrix}0&0&0\cr 0&0&-A_{2}^{T}\cr 0&-A_{2}&0\end{pmatrix}=\mathcal{M}_{3}-\mathcal{N}_{3},$ (8)
and proved that the stationary iteration methods corresponding to the above splittings are unconditionally convergent. They applied the matrices $\mathcal{M}_{i}$, $i=1,2,3$ as preconditioners to accelerate the convergence of GMRES for solving the system (9). In this paper, we propose a parameterized block splitting for another formulation of the equation (3) used in [15], investigate convergence of the method and induced preconditioner.
Throughout this paper, $\rho(A)$ and $\|A\|_{2}$ denote the spectral radius and the Euclidean norm, respectively. For a matrix $A\in\mathbb{R}^{r\times s}$, $A^{T}$ is used for the transpose of $A$. For two vectors $x$ and $y$, the notation $(x;y)$ is used for $(x^{T},y^{T})^{T}$. The null space of the matrix $A\in\mathbb{R}^{m\times n}$ is defined as $\mathcal{N}(A)=\{x\in\mathbb{R}^{n}:Ax=0\}$.
The remainder of this paper is organized as follows. In Section 2 we introduce the new iteration method and study its convergence properties. In Section 3, properties of the induced preconditioner are analysed. Some numerical results are given in Section 4. Finally, a short conclusion is presented in Section 5.
2 The proposed iteration method
We consider the following reformulation of the problem (3) used in [15]
$\begin{pmatrix}A_{1}&0&I\cr A_{2}&I&0\cr 0&-A_{2}^{T}&A_{1}^{T}\end{pmatrix}\begin{pmatrix}x\cr\delta_{2}\cr\delta_{1}\end{pmatrix}=\begin{pmatrix}b_{1}\cr b_{2}\cr 0\end{pmatrix},$ (9)
where $I$ denotes the identity matrix of the appropriate dimensions and $\delta=(\delta_{1};\delta_{2})=b-Ax$. It is noted that the coefficient matrix of this system is $\textbf{n}=2n+q$. By premultiplying the first row of the above system and letting $\hat{\delta}_{1}=A^{T}\delta_{1}$ and $\hat{b}_{1}=A_{1}^{T}b$, we obtain the following system
$\mathcal{A}\mathbf{x}\equiv\begin{pmatrix}P&0&I\cr A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr\delta_{2}\cr\hat{\delta}_{1}\end{pmatrix}=\begin{pmatrix}\hat{b}_{1}\cr b_{2}\cr 0\end{pmatrix}\equiv\mathbf{b},$ (10)
where $P=A_{1}^{T}A_{1}$. In this case, all the block diagonal matrices are nonsingular. We now consider the stationary iterative method induced by the parametrized block splitting (PBS) $\mathcal{A}=\mathcal{M}_{\alpha}-\mathcal{N}_{\alpha}$, where
$\mathcal{M}_{\alpha}=\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix},\quad\mathcal{N}_{\alpha}=\begin{pmatrix}0&0&-I\cr(\alpha-1)A_{2}&0&0\cr 0&0&0\end{pmatrix},$ (11)
and $\alpha$ is a positive number. The above splitting naturally introduces the stationary iteration method
$\mathbf{x}^{(k+1)}=\mathcal{G}_{\alpha}\mathbf{x}^{(k)}+\mathcal{M}_{\alpha}^{-1}\mathbf{b},$ (12)
for solving the system (10), where $\mathbf{x}^{(0)}$ is an initial guess and the matrix
$\mathcal{G}_{\alpha}=\mathcal{M}_{\alpha}^{-1}\mathcal{N}=\mathcal{I}-\mathcal{M}_{\alpha}^{-1}\mathcal{A},$
is the iteration matrix. Letting $\mathbf{x}^{(k)}=(x^{(k)};\delta_{2}^{(k)};\hat{\delta}_{1}^{(k)})$, the PBS iteration method can be rewritten as
$\begin{cases}x^{(k+1)}=-P^{-1}(\hat{\delta}_{1}^{(k)}+\hat{b}_{1}),\\ \delta_{2}^{(k+1)}=-\alpha A_{2}x^{(k+1)}+(\alpha-1)A_{2}x^{(k)}+b_{2},\qquad k=0,1,\ldots,\\ \hat{\delta}_{1}^{(k+1)}=A_{2}^{T}\delta_{2}^{(k+1)},\end{cases}$ (13)
where $\mathbf{x}^{(0)}=(x^{(0)};\delta_{2}^{(0)};\hat{\delta}_{1}^{(0)})$ is an initial guess.
The iterative scheme (12) converges to the solution of the system (10) for any arbitrary initial guess if and only if $\rho(\mathcal{G}_{\alpha})<1$ [13]. In the sequel we investigate the convergence of the proposed method. We first recall two lemmas.
Lemma 1.
[1] For two real numbers $b$ and $c$, the roots of the quadratic equation $x^{2}-bx+c=0$ are less than one in modulus if and only if
$|c|<1,\quad|b|<1+c.$
Lemma 2.
Let $A\in\mathbb{R}^{m\times n}$ and the matrix $A^{T}JA$ be SPD, where the matrix $J$ was defined in Eq. (2). Suppose that the matrix $A_{1}$ defined in Eq. (4) is of full column rank. If $\mu$ is an eigenvalue of the matrix $Q=(A_{1}^{T}A_{1})^{-1}(A_{2}^{T}A_{2})$, then $10\leq\mu<$ 1.
Proof.
It follows from Lemma 3.2 in [9] or Theorem 7.7.3 in [8]. ∎
Theorem 1.
Assume that the conditions of Lemma 2 hold. Then, the PBS iteration method converges to the solution of the system (10) for any arbitrary initial guess, i.e., $\rho(\mathcal{G}_{\alpha})<1$, if and only if
$0<\alpha<1+\frac{1}{\mu_{\max}}.$
where $\mu_{max}$ is the largest eigenvalue of $Q=(A_{1}^{T}A_{1})^{-1}(A_{2}^{T}A_{2})$.
Proof.
Let $(\lambda,\textbf{x})$ be an eigenpair of the iteration matrix $\mathcal{G}_{\alpha}$, i.e., $\mathcal{G}_{\alpha}\textbf{x}=\lambda\textbf{x}$. This means that $\mathcal{N}_{\alpha}\textbf{x}=\lambda\mathcal{M}_{\alpha}\textbf{x}$. So if we assume that $\textbf{x}=(x;y;z)$, then
$\begin{cases}-z=\lambda Px,\\ (\alpha-1-\alpha\lambda)A_{2}x=\lambda y,\\ z=A_{2}^{T}y.\end{cases}$ (14)
Premultiplying the second equation of (14) by $A_{2}^{T}$ and using the first equation, we get
$(\alpha-1-\alpha\lambda)A_{2}^{T}A_{2}x=\lambda A_{2}^{T}y=\lambda z.$
Now, substituting $z=-\lambda Px$ from the first equation in (14) in the above equation yields
$(\alpha-1-\alpha\lambda)P^{-1}A_{2}^{T}A_{2}x=-\lambda^{2}x.$
So, we deduce that
$(\alpha-1-\alpha\lambda)\mu=-\lambda^{2},$
where $\mu$ is an eigenvalue of $P^{-1}A_{2}^{T}A_{2}$. This equation can be written as
$\lambda^{2}-\alpha\mu\lambda+(\alpha-1)\mu=0.$ (15)
Now, from Lemma 1 we see that the method is convergent if and only if
$|(\alpha-1)\mu|<1,\quad\textrm{and}\quad\alpha\mu<1+(\alpha-1)\mu.$
From Lemma 2, we have $10\leq\mu<$ 1. For $\mu=0$, both of the above inequalities hold true. So, we assume that $\mu\neq 0$. Clearly, the second equation holds true. From the first equation we see that
$0<\alpha<1+\frac{1}{\mu},$
which gives the desired results. ∎
Remark 1.
From Theorem 1 we see the $\alpha=1$ is in the convergence interval of the PBS method. Hence, the method always is convergent for $\alpha=1$.
Theorem 2.
Assume that the conditions of Lemma 2 hold. The optimal value of the PBS method is given by
$\alpha_{opt}={\rm arg}\min_{0<\alpha<1+\frac{1}{\mu_{\max}}}\rho(\mathcal{G}_{\alpha})=\frac{2}{1+\sqrt{1-\mu_{\max}}}.$
and the corresponding optimal convergence factor is
$\rho(\mathcal{G}_{\alpha_{opt}})=\frac{\mu_{\max}}{1+\sqrt{1-\mu_{\max}}}.$
Proof.
We consider the equation (15) and without lose of generality we assume that $\mu=0$. In this case, this equation has two roots
$\lambda_{1}=\frac{\alpha\mu}{2}+\frac{\sqrt{\Delta}}{2}\quad\textrm{and}\quad\lambda_{2}=\frac{\alpha\mu}{2}-\frac{\sqrt{\Delta}}{2},$
where the discriminant is given by $\Delta=\alpha^{2}\mu^{2}-4(\alpha-1)\mu$.
We rewrite Eq. (15) in the form
$\frac{1}{\mu}\lambda^{2}=\alpha\lambda+(1-\alpha).$
Now, let
$f(\lambda)=\frac{1}{\mu}\lambda^{2},\quad\textrm{and}\quad g_{\alpha}(\lambda)=\alpha\lambda+(1-\alpha).$
For every $\alpha$, the line $g_{\alpha}$ passes through the point $(1,1)$, i.e., $g_{\alpha}(1)=1$ and the slope of $g_{\alpha}(\lambda)$ is $\alpha$. Figure 1 depicts the points of intersections of the functions $f$ and $g_{\alpha}$ for an arbitrary value of $\alpha$. Here, $\lambda_{1}$ and $\lambda_{2}$ are the roots of the quadratic equation (15). This figure shows that by increasing $\alpha,$ the value of $\max\{|\lambda_{1}|,|\lambda_{2}|\}$, decrease, while the line $g_{\alpha}$ gets tangent to $f(\lambda)$. In the tangent case, we have $\lambda_{1}=\lambda_{2}$ and it indicates that $\Delta=0$. With our assumption that $\mu\neq 0$, we obtain the quadratic equation
$\alpha^{2}\mu^{2}-4\alpha+4=0.$
This equation has two solutions
$\alpha_{\pm}=\frac{2}{1\pm\sqrt{1-\mu}}.$
It is easy to see that $\alpha_{-}\geq 1+1/\mu_{\max}$, and so $\alpha_{-}$ does not belong to the convergence interval. On the other hand, for $\alpha=\alpha_{+}$ we have
$\eta:=|\lambda_{1}|=|\lambda_{2}|=\frac{\alpha_{+}\mu}{2}=\sqrt{(\alpha_{+}-1)\mu}.$
For every $\alpha>\alpha_{+}$, the quadratic equation (15) has two complex roots
$\lambda_{1,2}=\frac{\alpha\mu}{2}\pm i\frac{\sqrt{-\Delta}}{2}$
where $i$ is the imaginary unit. It is not difficult to see that in this case
$|\lambda_{1}|=|\lambda_{2}|=\sqrt{(\alpha-1)\mu}>\eta.$
So, for fixed $\mu$, the best choice for $\alpha$ is
$\alpha_{+}=\frac{2}{1+\sqrt{1-\mu}}.$
Summarizing the above discussion the desired results are obtained.
Figure 1: Depiction of the functions $f(\lambda)=\frac{1}{\mu}\lambda^{2}$ and $g_{\alpha}(\lambda)=\alpha\lambda+(1-\alpha)$ in Theorem 2.
3 The PBS preconditioner
In the previous section, the PBS method is convergent if and only if $(0,1+1/\mu_{\max})$. It means that the eigenvalues of the matrix
$\mathcal{M}_{\alpha}^{-1}\mathcal{A}=\mathcal{I}-\mathcal{G}_{\alpha},$
are clustered in a circle with cener $(1,0)$ and radius 1. In this case, a Krylov subspace method like GMRES [14], would be well-suited for solving the preconditioned system [2]
$\mathcal{M}_{\alpha}^{-1}\mathcal{A}\textbf{x}=\mathcal{M}_{\alpha}^{-1}\mathbf{b}.$ (16)
In each iteration of a Krylov subspace method like GMRES it is needed to solve a system of equations
$\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}z_{1}\cr z_{2}\cr z_{3}\end{pmatrix}=\begin{pmatrix}w_{1}\cr w_{2}\cr w_{3}\end{pmatrix},$
can be done via the following steps
  • 1.
    Solve $Pz_{1}=w_{1}$ for $z_{1}$;
  • 2.
    Compute $z_{2}=w_{2}-\alpha A_{2}z_{1}$;
  • 3.
    Compute $z_{3}=w_{3}+A_{2}^{T}z_{2}$.
The coefficient matrix of the system in Step 1 is SPD. So, it can be solved directly using the Cholesky factorization for inexactly using the conjugate gradient method. The next theorem further analyses the eigenpairs distribution of the preconditioned matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$.
Theorem 3.
Assume that the conditions of Lemma 2 hold. Then, the eigenpairs of the matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$ are described as following
  • (i)
    If $\lambda\neq 1$ (for every $\alpha$), then the corresponding eigenvector is given by
    $\left(\frac{1}{\lambda-1}P^{-1}A_{2}^{T}y;y;A_{2}^{T}y\right),$
    where $y$ is nonzero vector.
  • (ii)
    If $\alpha=\lambda=1$, then every vector of the form $(x;y;0)$ with at least one of the vectors $x$ and $y$ being nonzero is the corresponding eigenvector.
  • (iii)
    If $\alpha\neq 1$ and $\lambda=1$, then every nonzero vector $(x;y;0)$ with $x\in\mathcal{N}(A_{2})$ is the corresponding eigenvector.
Proof.
Let $(\lambda,\textbf{x})$ be an eigenpair of the matrix $\mathcal{M}_{\alpha}^{-1}\mathcal{A}$, i.e., $\mathcal{M}_{\alpha}^{-1}\mathcal{A}\textbf{x}=\lambda\textbf{x}$. Letting $\textbf{x}=(x;y;z)$, this can be written as
$\begin{pmatrix}P&0&I\cr A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix}=\lambda\begin{pmatrix}P&0&0\cr\alpha A_{2}&I&0\cr 0&-A_{2}^{T}&I\end{pmatrix}\begin{pmatrix}x\cr y\cr z\end{pmatrix},$
which is equivalent to
$\begin{cases}(\lambda-1)Px=z,\\ (1-\alpha\lambda)A_{2}x=(\lambda-1)y,\\ (\lambda-1)(z-A_{2}^{T}y)=0.\end{cases}$ (17)
To prove $(i)$, we first show that $y\neq 0$. If $y=0$, then from the third equation in (17) we get $z=0$. Substituting $z=0$ in the first equation yields $x=0$. So, in this case we deduce that $\textbf{x}=(x;y;z)=0$ which is a contradiction with the vector x being an eigenvector. Now, from the third equation we see that $z=A_{2}^{T}y$. Substituting $z=A_{2}^{T}y$ in the first equation
$x=\frac{1}{\lambda-1}P^{-1}z=\frac{1}{\lambda-1}P^{-1}A_{2}^{T}y.$
To prove $(ii)$, if $\alpha=\lambda=1$ then it follows from (17) that $z=0$. It is noted that for every vectors $x$ and $y$ the Eq. (17) holds true that proves $(ii)$.
To prove $(iii)$, we see that $z=0$ and $(1-\alpha)A_{2}x=0$. We conclude that $A_{2}x=0$, meaning that $x\in\mathcal{N}(A_{2})$. Hence the desired results is obtained. ∎
4 Numerical experiments
I this section we present some numerical experiments to verify the theoretical results and show the effectiveness of the PBS preconditioner. Comparison with other recently proposed preconditioners are also given. All the numerical experiments have been done by some Matlab codes with a Laptop with 2.40 GHz central processing unit (Intel(R) Core(TM) i7-5500), 8 GB memory and Windows 10 operating system.
Example 1.
We consider the matrices $A_{1}$ and $A_{2}$ as following
$A_{1}=\begin{pmatrix}6&1&1\\ 2&4&5\\ 1&1&5\end{pmatrix},\quad A_{2}=\begin{pmatrix}2&1&1\\ 1&1&1\\ 1&2&2\\ 0&1&1\end{pmatrix},$
and $b_{1}$ and $b_{2}$ are two vectors whose entries are all ones. We have
$A_{1}^{T}A_{1}-A_{2}^{T}A_{2}=\begin{pmatrix}35&10&16\\ 10&11&19\\ 16&19&44\\ \end{pmatrix},$
which is an SPD matrix. So all the conditions of Theorems 1 and 2 hold true. On the other hand, we have $\mu_{\max}=0.4976$. So, the method is convergent for any $\alpha$ with
$3.0090<\alpha<1+\frac{1}{\mu_{\max}}=1+\frac{1}{0.4976}=$ 3.009 .
and optimal value for $\alpha$ is
$\alpha_{opt}=\frac{2}{1+\sqrt{1-0.4976}}=1.1704.$
Also, we have
$\rho(\mathcal{G}_{\alpha_{opt}})=\frac{0.4976}{1+\sqrt{1-0.4976}}=0.2912.$
We have solved the system (10) using the PBS method (Eq. (14)). The initial guess is set to be a zeros vector and the iteration is stopped as soon as residual norm of the system is reduced by a factor of $10^{11}$. The number of iterations for the convergence for the optimum value $\alpha$ and some other values have been presented in Table 1. The numerical results confirm the theoretical results.
Table 1: Numerical results for Example 1.
$\alpha$   0.7    0.8    1 $\alpha_{opt}=1.1704$    1.4    1.6    1.8
Iters 48 44 36 24 32 42 53
Example 2.
In this example, we use four matrices TOLS340,TOLS1090, TOLS2000 and TOLS400 from Matrix Market (https://math.nist.gov/MatrixMarket/). The Tolosa matrices are encountered in the stability analysis of an aircraft flight model. The notable modes of this system are characterized by complex eigenvalues, with their imaginary parts situated within a specified frequency range. The objective is to determine the eigenvalues that possess the largest imaginary components. This problem has been studied at CERFACS (Centre Européen de Recherche et de Formation Avancée en Calcul Scientifique) in collaboration with the Aerospatiale Aircraft division. These matrices with their generic properties, including the size ($n$), the number of nonzero entries ($nnz$) and the condition number estimation (Cond), are presented in Table 2.
Table 2: Matrix properties in Example 2.
matrix     $n$      $nnz$      Cond
TOLS 340 340 2196 $052.4e+$ 05
TOLS1090 1090 3546 $062.1e+$ 06
TOLS2000 2000 5184 $066.9e+$ 06
TOLS4000 4000 8784 $072.7e+$ 07
We set $A_{1}=$ TOLS 340, TOLS1090, TOLS2000 and TOLS4000. In this case, we have $p=n=340,1090,2000$ and $4000$. For each system we set $q=n$ and $A_{2}=6I_{n}$, where $I_{n}$ is the identity matrix of order $n$. The entries of the vectors $b_{1}$ and $b_{2}$ are set to be all ones. We solve the system (10) with restarted version of GMRES(10) in conjunction with the PBS preconditioner. Left-preconditioning is used. A zero vector is used as an initial guess and the iteration is stopped as soon as the residual of the original system is reduced by a factor of $10^{11}$. The parameter $\alpha$ in the PBS method is set to be $1$. We also compare the numerical results of PBS preconditioner with the preconditioners induced by the block splittings (6), (7) and (8) (denoted by BS1, BS2 and BS3) applied to the system (10). Numerical results have been presented in Table 2. In this table the number of iterations (Iter), elapsed CPU time (CPU),
$Rel=\frac{\|\textbf{r}_{k}\|_{2}}{\|\textbf{r}_{0}\|_{2}},$
with $r_{k}$ is the residual vector at iteration $k$,
$Err=\frac{\|\textbf{x}_{k}-\textbf{x}^{*}\|_{2}}{\|\textbf{x}^{*}\|_{2}},$
[…]
Table 3: Numerical results for Example 2.
   $m\times n$ PBS BS1 BS2 BS3 No-Prec
[…]
CPU 0.08 0.10 0.09 0.10
RES 9.13e-12 3.79e-12 3.52e-13 7.31e-12
ERR 9.22e-12 3.07e-11 9.16e-13 9.06e-12
[…]
CPU 0.09 0.11 0.10 0.11
RES 2.58e-12 2.71e-12 6.05e-12 6.26e-12
ERR 3.61e-12 1.83e-11 1.76e-11 7.87e-12
[…]
CPU 0.09 0.12 0.10 0.13
RES 1.91e-12 4.12e-12 4.42e-12 4.78e-12
ERR 3.02e-12 3.45e-11 1.28e-11 5.05e-12
[…]
CPU 0.11 0.14 0.12 0.15
RES 7.96e-12 8.98e-12 3.04e-12 3.07e-12
ERR 1.81e-11 1.44e-10 9.70e-12 5.23e-12
Example 3.
We consider the partial differential equations
$\left\{\begin{array}[]{rll}-\Delta u+\sin(x+y)\frac{\partial u}{\partial x}+\cos(x-y)\frac{\partial u}{\partial y}+50(x+y)u=f,&{\rm in}&\Omega=[0,1]\times[0,1],\\ u=g,&{\rm on}&\partial\Omega,\end{array}\right.$ (18)
where the function $g$ is known.The discretization of this equation employing the second-order central difference scheme on an $(n_{0}+2)$ grid in all directions of $\Omega$, with a mesh size of $h=1/(n_{0}+1)$, yields a system of linear equations with the coefficient matrix $A_{1}\in\mathbb{R}^{n\times n},$ in which $n=n_{0}^{2}$. We set $A_{2}=0.7I_{n}$, where $I_{n}$ is the identity matrix of order $n$.
We set $n_{0}=85,90$ and $95$. Consequently, the resulting matrices $A_{1}$ are of orders $7225$, $8100$, and $9025$, respectively. In this case, the order of matrices $A$ is $14450\times 7225$, $16200\times 8100$ and $18050\times 9025$. All the assumptions are as the previous example, except here we have used the complete version of GMRES. All the other assumptions are the same as in the previous example. Numerical results are presented in Table 3. As we observe the PBS preconditioner outperforms the other preconditioners.
Table 4: Numerical results for Example 3.
   $m\times n$ PBS BS1 BS2 BS3 No-Prec
[…]
CPU 0.10 0.18 0.11 0.20
RES 7.49e-12 5.86e-12 4.63e-13 1.47e-12
ERR 4.30e-09 1.01e-09 1.68e-09 1.68e-09
[…]
CPU 0.11 0.21 0.13 0.23
RES 7.52e-12 5.87e-12 5.22e-13 1.65e-12
ERR 3.43e-09 1.88e-09 1.88e-09 1.88e-09
[…]
CPU 0.13 0.24 0.15 0.27
RES 7.53e-12 5.88e-12 5.71e-13 1.75e-12
ERR 5.85e-09 1.65e-09 1.64e-09 1.65e-9
5 Conclusions
We have presented a parameterized block splitting iteration method for solving indefinite least squares problem. We have proved that the iteration method is convergent under a condition. Optimal value of the parameter has been computed. Eigenpair properties of the preconditioned matrix has been studied. We have compared the numerical results of the preconditioner induced by the proposed method with three recently presented preconditioners. The numerical results show that the proposed preconditioner in more efficient that others.
Conflicts of interest
This work does not have any conflicts of interest.